# GlucOS: Security, correctness, and simplicity for automated insulin delivery

### v3

@@ 9.1.1. Without algorithmic security mechanism
Predictive ML model baseline With the predictive ML model, we followed all ML best practices in carefully tuning hyperparameters and evaluating the model on simulation scenarios that were different from those used in training. Figure 6 shows the average proportion of time spent in the safe target range (TIR) by virtual humans from different age groups when employing the predictive ML model. We see that individuals across all groups spend over 85% of their time in range compared to 77% when using only the reactive safe model. These results provide a baseline for an algorithm that is effective in managing T1D.
@@
Incorrect ML dosing Since GlucOS is designed to be flexible with any predictive algorithm, it could be subject to attacks from malicious ML algorithms. In this section, we discuss an experiment with a malicious algorithm that intentionally doses too much insulin. Concretely, our malicious algorithm doses ten times the amount of insulin computed by the previously described predictive ML model. We ran this algorithm within the simulator on the same virtual human on the same scenarios as the previous experiment.
Figure 7 shows the CGM readings on three randomly chosen virtual humans on a particular scenario when we dose more. The CGM readings of all 3 individuals fully go below the hypogycemic bound, which would have killed them in the real-world.
@@
Across all virtual humans, we see an average drop in TIR from 88.44% to 0% when dosing too much insulin. Correspondingly, we observe an average increase in the time spent in hypoglycemia from 0% to 100%. We also observed similar results with experiments where we intentionally underdosed insulin. In these experiments, all virtual humans spend most of their time above the hyperglycemic bound which wouldn’t have been fatal but would have led to severe health consequences in the real world. These results motivate the need for our algorithmic security mechanism.
@@ 9.1.2. With algorithmic security mechanism
Impact on incorrect dosing. We repeat experiments from the previous section with the same scenarios. We run the same malicous model to inject too much insulin, but also run our reactive safe model (Section 3) to bound insulin injection.
Figure 8 show CGM readings on the same three virtual humans on the same scenarios as the experiment in the previous section. We see that all of them spend over 70% of their time in range, showing that GlucOS protects individuals even in presence of malicious overdosing. On average, across all virtual humans, we only see a drop in TIR from 88.44% with correct dosing to 78.22% when employing safety while overdosing. We also report similar results on experiments with a malicious algorithm that underdoses insulin.
While the times spent in range are less than those of correct dosing, the impact of incorrect dosing is significantly dampened with individuals spending the recommended amount of time in range, despite facing active attacks.
@@
Comparing data in Figure 6 we see that there is only a slight drop in the average proportion of time spent in range by virtual humans across all age groups when employing GlucOS’s algorithmic security mechanism. When employing the reactive safe model as is to calculate insulin doses, we report an average TIR of 82.48% which is lower than the average TIR of 86.84% observed when running the predictive ML model with GlucOS’s security mechanism. This shows that GlucOS’s algorithmic security guarantees do not come at the cost of effective T1D management from predictive ML.
9.2. Does GlucOS ensure pump operation while overcoming malicious pump drivers?
Our mechanism should ensure that all communications go through with a benign pump driver and should catch all manipulated communication with a malicious pump driver.
Impact on malicious pump drivers. We attached an Omnipod DASH insulin pump to a stuffed animal (Figure 9). We then modified the OmniBLE pump driver to intentionally add, drop and modify insulin commands. We detected all manipulated commands thereby showing that our driver security mechanism has 0 false negatives in detecting them.
Impact on benign pump drivers. For one week, we logged messages passed between the BioKernel, the unmodified OmniBLE pump driver and the pump on an individual running GlucOS (Section 10). We report 100% consistency between the commands sent to OmniBLE and commands issued from OmniBLE showing that GlucOS’s driver security mechanism does not incur any false positives.
@@ 10. Using GlucOS in the real world
In this section, we report on our experiences working with seven individuals using GlucOS to help manage their T1D. Six people are students who only used the predictive alerting feature of the BioKernel (Section 6) for two months, from June 2024 - August 2024. One individual, who we will call Bob, has been using the full GlucOS system since November 2023 for insulin delivery and predictive alerting.
@@ 10.1. Ethical considerations
In this section, we outline the steps we take to ensure that we uphold high ethical standards and ensure that the people using our system, are safe. Our user study for the six students was approved by our university’s IRB, and Bob’s use GlucOS was determined by our university’s IRB to be “self evaluation” and exempt from IRB.
One of the co-authors of this paper is a board-certified Endocrinologist, who specializes in T1D. They designed the safety protocol, which defines the specific criteria we use to stop the study, if needed. The Appendix will include a more detailed discussion of our ethics and safety protocols.
@@ 10.2. Is an ML-based closed-loop system safe and effective in practice?
Overall, Bob had the best results when using ML. His time spent in range increased from 96.3% to 97.2%, and notably he spent no time in hypoglycemia (below 70 mg/dl), down from 0.67% when not using ML.
Next, we measure how often the predictive ML model was responsible for programming the pump vs our reactive safe model. For the one week when Bob was using the predictive ML model, 33.0% of the time both the predictive ML model and reactive safe model issued the same commands, 39.9% of the time the BioKernel used the predictive ML model’s commands, and 27.1% of the time the BioKernel used the reactive safe model’s commands. We believe that this distribution represents a suitable balance for providing the ML with enough flexibility to improve outcomes, while having tight enough bounds to limit potential damage from bad or malicious ML predictions.
@@ 10.3. Do biological invariant violations trigger in practice?
These results suggest that although natural variations in glucose metabolism do lead to biological invariant violations, none of them were long enough to trigger our most severe action of transitioning to manual mode after two hours.
We reviewed one weeks worth of violations with Bob manually to determine the cause. During the most recent week of data, Bob had five sequences of 30 minutes or longer and all were caused by exercise.
@@ 10.4. Are predictive alerts useful for security?
To measure predictive alerts as a security mechanism (Section 6), we recruited six students who live with T1D and had them use the GlucOS predictive alerts to complement their default CGM alerts. Our primary metric is time spent below 70 mg/dl, where a decrease in the amount of time spent in hypoglycemia suggests that predictive alerts are effective to empower people to mitigate security attacks.
@@
Overall, our results suggest that predictive alerts can be a useful last line of defense for trustworthy automated insulin delivery systems.
@@ 10.5. How does the BioKernel perform?
We look at two aspects of system performance when evaluating our real-world use of GlucOS. First, we measure how long our closed-loop algorithm takes to run and second, we count how often the closed-loop algorithm runs successfully end-to-end. We use Bob’s iPhone 14 to collect all of the data we show here and these numbers are averages across the entire eleven month period.
@@ 11. Related work
In addition to the work already discussed in this paper, there are several other related works.
The complexity of automated insulin delivery systems and the fact that they make real-time healthcare decisions, makes their security paramount. Current research on security for these systems focuses primarily on evaluating algorithms using simulations (zhou2021data, ; zhou2023knowsafe, ). We argue that this approach falls short of addressing the full spectrum of challenges inherent in building trustworthy systems that operate within real-world constraints. Real-world systems must cope with runtime platform limitations, changes in human physiology, and the necessity of human interaction with the system. Plus, algorithms and simulations model attacks abstractly through their glucose response, missing out on details like which components of the system are likely to be compromised or how the system architecture can prevent damage from these potentially vulnerable components. Our experience shows that when faced with these practical challenges, algorithms designed in simulations often prove insufficient or inapplicable. These real-world complexities require a fundamentally different approach to designing security for automated insulin delivery systems. To address these challenges, GlucOS moves beyond simulation-based designs and tackles the challenges of real-world systems.
Previous research has looked at the security of implanted medical devices in general (halperin2008security, ; burleson2012design, ; rushanan2014sok, ), in addition to looking at insulin pumps in particular (li2011hijacking, ; paul2011review, ), with more recent work looking at providing improved security (marin2016feasibility, ; ahmad2018securing, ). Also, recent work has looked at applying formal methods to insulin pumps for high assurance (panda2021secure, ). These works focus on devices and their communication channels. In contrast, with GlucOS we assume that these devices are correct and secure and focus our efforts on the software we use to run the automated insulin delivery system.
@@ 12. Conclusion
First, our approach to algorithmic security used classic security techniques, like reference monitors and the Simplex architecture. However, it was our novel insight around all correct algorithms dosing the same amount of insulin that led us to these classic techniques. Plus, from our exploration, the most interesting part was not the ML, it was around the principles of reactive safe models and in particular our principle that these safety-critical algorithms should be reactive.
Second, our pump driver security went beyond traditional specification and checking. We needed to actively program the pump to handle dropped cancel commands, which was an unexpected but crucial insight.
@@
Third, the biological invariant emerged as a key check in our system, addressing a blind spot in typical simulations that ignore changing human dynamics. This forced us to develop novel formal methods to handle scenarios where our glucose metabolism models could be drastically wrong.
Fourth, our approach to human interaction in the system was nuanced. We found that simply alerting humans or suspending insulin delivery were not always appropriate. For security alerts, personalization and customization were important. Stopping insulin delivery could be dangerous if the user has high glucose and insufficient insulin on board, which required more context-aware responses.

### v4

@@ 9.1.1. Without algorithmic security mechanism
Predictive ML model baseline With the predictive ML model, we followed all ML best practices in carefully tuning hyperparameters and evaluating the model on simulation scenarios that were different from those used in training. Figure 5 shows the average proportion of time spent in the recommended healthy range (TIR) by virtual humans from different age groups when employing the predictive ML model. We see that individuals across all groups spend over 85% of their time in range compared to 77% when using only the reactive safe model. These results provide a baseline for a predictive algorithm that is effective in managing T1D.
@@
Incorrect ML dosing In this section, we discuss an experiment with a malicious algorithm that intentionally doses too much insulin. Concretely, our malicious algorithm doses ten times the amount of insulin computed by the previously described predictive ML model. We ran this algorithm within the simulator on the same virtual humans on the same scenarios as the previous experiment.
Figure 6 shows the CGM readings on three randomly chosen virtual humans. The readings fully go below the hypogycemic bound, which would have killed them in the real world.
@@
Across all virtual humans, we see an average drop in TIR from 88.44% to 0% when dosing too much insulin. Correspondingly, we observe an average increase in the time spent in hypoglycemia from 0% to 100%. We report similar results with intentional underdosing in Appendix C. These results demonstrate the need for algorithmic security.
@@ 9.1.2. With algorithmic security mechanism
Impact on incorrect dosing. We repeat experiments from the previous section with the same simulation scenarios. We run the same malicous model to overdose insulin, but also run our reactive safe model (Section 4).
Figure 7 show CGM readings on the same three virtual humans on the same scenarios as the experiment in the previous section. We see that all of them spend over 70% of their time in the healthy range, showing that GlucOS protects individuals from malicious overdosing. On average, across all virtual humans, we only see a drop in TIR from 88.44% with correct dosing to 78.22% when employing GlucOS while overdosing. We also report similar results on experiments with a malicious algorithm that underdoses insulin in Appendix C.
While the times spent in range are less than those of correct dosing, the impact of incorrect dosing is significantly dampened with individuals spending the recommended amount of time in the healthy range, despite active attack.
@@
Comparing data in Figure 5 we see that there is only a slight drop in the average proportion of time spent in range by virtual humans across all age groups when employing GlucOS’s algorithmic security mechanism. When employing the reactive safe model as is to calculate insulin doses, we report an average TIR of 82.48% which is lower than the average TIR of 86.84% observed when running the predictive ML model with GlucOS’s security mechanism. This shows that GlucOS’s algorithmic security mechanism does not impede effective T1D management.
@@ 10. Using GlucOS in the real world
In this section, we report on our experiences working with seven individuals using GlucOS to help manage their T1D. Six people are students who only used the predictive alerting feature of the BioKernel (Section 6) for two months, from June 2024 - August 2024. One individual, who we will call Bob, has been using the full GlucOS system since November 2023 for insulin delivery and predictive alerting. The number of participants in our study is roughly consistent with existing research in human factors in computing (glucomaker, ; moont1d, ; context_aware_t1d, ; t1d_children, ). More importantly, we report results on indiviuals using GlucOS to manage their diabetes in the real world.
@@ 10.1. Ethical considerations
In this section, we outline the steps we take to ensure that we uphold high ethical standards and ensure that the people using our system, are safe. Our user study for the six students was approved by our university’s IRB. Bob’s use GlucOS was determined by our university’s IRB to be “self evaluation” and exempt from IRB.
One of the co-authors of this paper is a board-certified Endocrinologist, who specializes in T1D. They designed the safety protocol, which defines the specific criteria we use to stop the study, if needed. Appendix A includes a more detailed discussion of our ethics and safety protocols.
@@ 10.2. Is an ML-based closed-loop system safe and effective in practice?
Overall, Bob had the best results when using ML. His time spent in range increased from 96.3% to 97.2%, and notably he spent no time in hypoglycemia (below 70 mg/dl), going down from 0.67% when not using ML. More importantly, Bob’s A1C, which captures his long-term glucose concentration, dropped to 5.8% as a result of using GlucOS. This result demonstrates a significant health outcome since it matches up to non-diabetics who have an A1C of 5.6% or lower.
Next, we measure how often the predictive ML model was responsible for programming the pump vs our reactive safe model. For the one week when Bob was using the predictive ML model, 33.0% of the time both the predictive ML model and reactive safe model issued the same commands, 39.9% of the time the BioKernel used the predictive ML model’s commands, and 27.1% of the time the BioKernel used the reactive safe model’s commands. The reactive safe model having to take over to issue commands is not surprising since ML models are not immune to mispredictions, especially when taking the variability of the real world into account. We believe that this distribution represents a suitable balance for providing the ML with enough flexibility to improve outcomes, while having tight enough bounds to limit potential damage from bad or malicious ML predictions.
As of July 2025, Bob has swapped the predictive ML model with a rule-based model similar to OpenAPS’ model described in Section 1.
@@ 10.3. Do biological invariant violations trigger in practice?
These results suggest that although natural variations in glucose metabolism do lead to biological invariant violations, none of them were long enough to trigger our most severe action of transitioning to manual insulin dosing after two hours. We manually reviewed one weeks worth of Bob’s data and discovered that all five violations within this period were caused by exercise.
@@ 10.4. Are predictive alerts useful for security?
To measure predictive alerts as a security mechanism to avert hypoglycemia (Section 6), we recruited six students who live with T1D and had them use the GlucOS predictive alerts to complement their default CGM alerts. Our primary metric is time spent below range (i.e., below 70 mg/dl), where a decrease in the amount of time spent in hypoglycemia suggests that predictive alerts are effective to empower people to mitigate the most immediate and dangerous security risks.
@@
We also report positive feedback from all participants during monthly interviews. Other than customization, participants acknowledged the importance of being able to reason about the alerts once we explained that we used linear regression for prediction. They reported that this helped forge trust in our system which in turn reduced their cognitive load with diabetes. All participants have continued using our system even after the study’s conclusion.
Overall, our results show how predictive alerts are useful for automated insulin delivery.
@@ 10.5. How does the BioKernel perform?
We evaluate two aspects of GlucOS’s system performance in the real world. First, we measure how long our closed-loop algorithm takes to run and second, we count how often the closed-loop algorithm runs successfully end-to-end. We use Bob’s iPhone 14 to collect all of the data we show and these numbers are averages across the entire eleven month period.
@@ 11. Related work
The complexity of automated insulin delivery systems and the fact that they make real-time healthcare decisions, makes their security paramount. Current research on security for these systems focuses primarily on evaluating algorithms using simulations (zhou2021data, ; zhou2023knowsafe, ). Our experience shows that when faced with the practical challenges of real systems, algorithms designed in simulations often prove insufficient or inapplicable. These real-world complexities require a fundamentally different approach to designing security for automated insulin delivery systems. GlucOS moves beyond simulation and tackles the challenges of real-world systems.
Previous research has looked at the security of implanted medical devices in general (halperin2008security, ; burleson2012design, ; rushanan2014sok, ), in addition to looking at insulin pumps in particular (li2011hijacking, ; paul2011review, ), with more recent work looking at providing improved security (marin2016feasibility, ; ahmad2018securing, ). Also, recent work has looked at applying formal methods to insulin pumps for high assurance (panda2021secure, ). They focus on devices and their communication channels. With GlucOS, we assume that these devices are secure and focus on automated insulin delivery software.
@@ 12. Conclusion
First, our approach to algorithmic security used classic security techniques, like reference monitors and the Simplex architecture. However, our novel insight around all algorithms dosing the same amount of insulin led to the application of these classic techniques. Plus, from our exploration, the most interesting part was not the use ML, but the principle of repurposing reactive models for security.
@@
Second, the biological invariant emerged as a key check in our system, addressing a blind spot in typical simulations that ignore changing human dynamics. This forced us to develop novel formal methods to handle scenarios where our glucose metabolism models could be drastically wrong.
Third, our approach to human interaction in the system was nuanced. We found that simply alerting humans was not sufficient. For security alerts, personalization, customization, and simplicity was important.